An MTN cellular-phone user notices that her cellular phone accounts for three consecutive months were significantly different. In May, her account was one and a half times more than her account in June. Her account in July was R 50 more than her account in June. In total, she spent R575 in cellular-phone accounts in the three months. What was her account for each month?
1
Expert's answer
2019-04-08T12:26:49-0400

Let's consider account in June as x. So account in May was 1.5x as well as (50+x) in July.

"1.5x+x+50+x=575"

"3.5x=525"

"x= \frac{525}{3.5}"

"x=150"


"1.5x=1.5 \cdot 150 = 225"

"50+x=50+150=200"


Answer: account in May was R225; account in June was R150; account in July was R200.